OpenStudy (anonymous):

Write an equation for a graph that is the set of all points in the plane that are equidistant from the point F(2,0) and the line x=-2

OpenStudy (anonymous):

\[ (x+2)^2=(x-2)^2 + (y-0)^2\\ \]

OpenStudy (anonymous):

\[ x^2 + 4 x + 4 = x^2 - 4 x + 4 + y^2\\ 8x = y^2 \]